State whether the following statements are true or false. Give reasons for your answers 3) The domain of the f/g where f(x, y) =2xy and g(x, y) =x^2+y^2 is R^2.
1
Expert's answer
2020-03-23T13:55:52-0400

The given function are ,

"f(x,y)=2xy \ and \ g(x,y)=x^2+y^2"

We have to check whether, "\R^2" is a domain of "\frac{f}{g}" Or not?

Clearly,

"f:\R^2\rightarrow \R" and "g:\R^2\rightarrow \R" and "g(x,y)=0" "\iff" "x=y=0"

Hence ,Domain of "\frac{f}{g}" is "\R^2 -(0,0)" .

Hence the statement is false.